Use simpson's method to approximate

8

∫ (x^2-x+3)dx with 8 sub-interval



1
Expert's answer
2021-06-02T09:18:24-0400

Approximate the integral


"\displaystyle\int_{0}^{8}(x^2-x+3)dx"

with "n=8" using the Simpson's rule.


"\displaystyle\int_{a}^{b}f(x)dx\approx\dfrac{\Delta x}{3}(f(x_0)+4f(x_1)+2f(x_2)+..."

"+2f(x_{n-2})+4f(x_{n-1}+f(x_n))"

where "\Delta x=\dfrac{b-a}{n}"

We have that "a=0, b=8, n=8." Therefore "\Delta x=\dfrac{8-0}{8}=1."

Divide the interval"[0, 8]"into "n=8" subintervals of the length "\Delta x=1" with the following endpoints: "a=0, 1, 2, 3, 4, 5, 6, 7, 8=b."

Now, just evaluate the function at these endpoints.


"f(x_0)=f(0)=3"

"4f(x_1)=4f(1)=12"

"2f(x_2)=2f(2)=10"

"4f(x_3)=4f(3)=36"

"2f(x_4)=2f(4)=30"

"4f(x_5)=4f(5)=92"

"2f(x_6)=2f(6)=66"

"4f(x_7)=4f(7)=180"

"f(x_8)=f(8)=59"


"\dfrac{1}{3}(3+12+10+36+30+92+66+180+59)"

"=\dfrac{488}{3}\approx162.6667"

"\displaystyle\int_{0}^{8}(x^2-x+3)dx\approx\dfrac{488}{3}\approx162.6667"